Jeff is three times as old as his son. Jeff will be twice as old as his son in 13 years. How old is Jeff now
Answer by JBarnum(2146)

Jeff Son
J=3S
J+13=2(S+13)
J+13=2S+26
J=2S+13
3S=2S+13
S=13
J=3(13)
J=39